OpenStudy (anonymous):

You have 4/6 cup of oatmeal in your cupboard. A recipe for cookies calls for 1/4 cup of oatmeal. How much oatmeal would you have left if you made the cookies?

OpenStudy (anonymous):

\[\frac{4}{6}-\frac{1}{4}\]

OpenStudy (anonymous):

btw \(\frac{4}{6}=\frac{2}{3}\) so this is the same as \[\frac{2}{3}-\frac{1}{4}\]

OpenStudy (anonymous):

you know how to do that subtraction?

OpenStudy (anonymous):

don't the denominators have to be equal numbers

OpenStudy (anonymous):

make them equal by writing \[\frac{2}{3}=\frac{8}{12}\] and \[\frac{1}{4}=\frac{3}{12}\]now you have \[\frac{8}{12}-\frac{3}{12}\]

OpenStudy (anonymous):

so then the answer is 5/12 ? im terrible at math sorry lol

OpenStudy (anonymous):

or you can always subtract via \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bc}\] so you would get \[\frac{2}{3}-\frac{1}{4}=\frac{2\times 4-3\times 1}{3\times 4}\]

OpenStudy (anonymous):

and yes, the answer is \(\frac{5}{12}\)

OpenStudy (anonymous):

got another?